For the following reaction at 373 K, Kc = 0.50. If the initial concentration of N2O4 is 0.152 M, what is the equilibrium concentration of NO2?



N2O4 (g) ⇆ NO2 (g)

1
Expert's answer
2022-04-25T17:04:04-0400

Solution:

Balanced chemical equation:

N2O4(g) ⇆ 2NO2(g)

For the given reaction,



ICE Table:



According to the ICE Table:

[NO2] = 2x

[N2O4] = 0.152 − x

Kc = 0.50

Therefore,

0.50 = (2x)2 / (0.152 − x) = (4x2) / (0.152 − x)

4x2 + 0.50x − 0.076 = 0

Solving this quadratic equation, we get:

x = 0.089


[NO2] = 2x = 2 × 0.089 = 0.178 M

[NO2] = 0.178 M


Answer: The equilibrium concentration of NO2 is 0.178 M